# Examining the Prevalence and Dynamics of AI-Generated Media in Art Subreddits

### v1

@@ Introduction
Having gathered a dataset of potentially AI-relevant image-based posts based on keyword match, we further classified the posts using large language models (LLMs) (Abburi et al. 2023; Zhang et al. 2024; Pangakis and Wolken 2024) First, since keyword matching can be imprecise, we use classifiers to filter the dataset to only retain posts and comments genuinely pertaining to AI-generated images. Second, we differentiate between text that identifies the image in the post as being AI-generated, versus those that broadly mention AI-generated images.
@@
To this end, we developed textual prompts and utilized the ChatGPT API [^9] https://platform.openai.com/docs/api-reference/introduction. Three distinct prompts were developed to correspond with each of the three type of text in our dataset: post’s title or body, comments made by the author of the post, and comments by other users (see Appendix C). These prompts were designed following the principles of chain-of-thought prompting (Wei et al. 2022). Each prompt required the language model to provide reasoning before giving a classification decision. Additionally, each prompt included nine examples from the dataset, implementing an in-context few-shot learning approach (Wei et al. 2022).
@@
We evaluated the accuracy of each of the three prompts using three validation sets. We randomly sampled 100 texts from each type, and manually labeled them according to the established coding. That is, the texts were labeled as follows: (A) Implies AI involvement, when the text explicitly or implicitly suggests that AI tools were involved in generating the attached image; (B) Does not imply AI, when the text mentions AI-generated images, but does not imply that the attached image is AI-generated (including texts that suggests that the attached image is not AI-generated); or (C) Unrelated to AI, only when the text is completely unrelated to AI-generated images (regardless of whether the text refers to the image in the post)—in most cases, a false positive of the keyword match. Next, we classified the three validation sets using the corresponding prompts. This classification task achieved an accuracy of 93% for the title of the posts, 82% for comments made by the author of the post, and 84% for comments from other accounts.
@@
We classified all data entities in the dataset using these prompts, following a process inspired by the QuaLLM framework (Rao et al. 2024). According to this process we submitted the prompt, consisting in each case of the relevant class explanation and the in-context examples, with every 100 input texts at a time [^10] Note that during the evaluation phase we also compared the accuracy values obtained by submitting different numbers of entities as input (i.e., 10 instead of 100) and achieved identical results.. Similarly, we requested that the LLM return the final output as a dictionary, including both the reasoning and classifications for all these 100 texts.
@@
In this paper, we focus mainly on texts, either in posts or comments, that imply that the image attached to the post was AI-generated (the first category of labels above). First, looking at people who make image posts in these communities, we identify on posts where the author implied or admitted that the image was AI-generated, whether in the title, or in one of the comments. We refer to posts where the creator itself implied the image was AI generated as “Transparent AI posts”. There are 5,998 distinct posts like these where the creator provides this information in the title (5,144 times) and/or comments (1,243). Second, we are interested in posts where other users or accounts make comments that refer in some way to the posted image as AI-generated. We refer to these posts where other accounts highlighted AI use as “Suspected AI posts”. There were a total of 7,322 comments from other accounts claiming AI was used, made on 4,375 distinct Suspected AI posts. The overlap between the two sets of posts was small—576 posts where the author transparently posted AI and commenters specifically referred to AI generation in their comments.
Categorization of the Communities
@@
We categorized the 57 subreddits into two types of communities: those with rules that explicitly ban AI images, and those whose rules do not mention AI. Previous papers have explored the reactions of community moderators to AI-generated content, including focusing on the decisions of Reddit moderators to enact rules restricting the use of AI, as well as the responses of their communities (Lloyd, Reagle, and Naaman 2023; Shankar and Sim 2024). We set out to distinguish the different communities based on their expressed approach to AI content. Using the data provided via the PRAW API [^11] https://praw.readthedocs.io/en/latest/index.html, we examined each subreddit’s API description, public description, and rules. If any of these fields indicated that AI-generated content (or particularly AI-generated images or art) is not permitted, we categorized the subreddit as “AI-disallowed community”. Otherwise, if the publication of AI-generated content was not mentioned in the subreddit’s description or rules, it was categorized as “AI-neutral community” [^12] In our dataset, no subreddit explicitly allows AI art in their description or rules.. Among the 57 subreddits, 24 explicitly forbid AI, while the remaining 33 subreddits do not reference AI in their descriptions and rules.
@@
To obtain a more comprehensive understanding and baseline for the subreddits in each category, we utilized torrents (Watchful1 2024) to count the total number of posts that included a media item in these communities, regardless of whether they contained the AI-related keywords. We refer to these posts that included a media item, such as an image or video, as “all image-based posts”.
This quantitative comparison reveals several unexpected findings. Although the majority of posts of any kind within these subreddits are image-based, comprising 75% in AI-disallowed communities and 77% in AI-neutral communities, only a minimal percentage of the posts included texts implying AI involvement in creating the image. Specifically, AI-claiming texts were presented only in 0.21% of image-based posts in AI-disallowed communities, and 0.12% in AI-neutral communities, including texts made by the author of the post (referred to as Transparent AI posts) and those made by commenters (referred to as Suspected AI posts). Furthermore, when AI involvement was mentioned in relation to the creation of the image, it was mostly by the content creators themselves, accounting for 57% (4,458 out of 7,792) of cases in AI-disallowed communities, and 60% (1,540 out of 2,581) in AI-neutral communities. Finally, for 87% (2,894 out of 3,334) of the Suspected AI posts in the AI-disallowed communities, i.e. posts where a non-author commenter implied AI involvement, the comment served as the first indication that the attached image was AI-generated, without the author of the post disclosing the use of AI. The same percentage was observed in the second category as well. We explore these differences further in our analysis.
@@
Table 1: Characteristics of the AI-disallowed and AI-neutral communities.
Analyses and Findings
@@
In this section, we use the dataset we collected and refined in previous sections to examine variations in volume, participation, and feedback patterns across AI-disallowed communities and AI-neutral communities. This analysis highlights the distinctions between these two types of communities, as well as the differences in trends for the different types of posts — Transparent AI posts and Suspected AI posts.
Rise of AI Posts and Comments
@@
Figure 1: The volume of posts over time in AI-disallowed communities (red) and AI-neutral communities ( green). The x-axis of the figure represents time (month), and the y-axis represents the number of posts. Plots 1-2 (top two) show the total volume of all image-based posts; plots 3-4 show Transparent AI posts, and plots (5-6) show the number of Suspected AI posts.
@@
We conducted an analysis of post volume over time across different communities to understand how posting patterns shifted in response to the introduction and increasingly mainstream adoption of generative AI throughout 2022. Specifically, we examined temporal patterns of posts containing AI-generated images according to our classification, and compared these trends to the overall volume of image-based posts within the subreddits we examined. We explore how these trends vary over time between the AI-disallowed communities and the AI-neutral communities. Figure 1 shows the number of posts over time between both types of communities: AI-disallowed communities in red, and AI-neutral communities in green. The figure presents the number of posts for various data segments, enumerating from the top panel: (1-2) all image-based posts, (3-4) Transparent AI posts when authors admitted to posting AI-generated images, and (5-6) Suspected AI posts when commenters implied that the image in the post is AI-generated. For instance, Panel 2 shows that between March and August 2020, approximately 35,000 to 40,000 image-based posts were published in AI-neutral communities (which do not mention AI posting in their community rules). In contrast, two years later (March-August 2022), only approximately 25,000 to 30,000 image-based posts were published in the same subreddits.
@@
While the overall number of image-based posts has been on a decline since August 2021 in both community types (panels 1-2), the number of Transparent AI posts began to increase during the same period (panels 3-4). This difference suggests that the rise in generative AI posts is not due to broader trends in posting volume across subreddits, but indicates specific trends related to AI-generated images.
@@
Participation and Newcomers
The previous sections focused on volume of posts and discussions over time and in relation to other image posts in the same communities. In this section, we examine the pattern of participation in creating AI-related posts across the different art communities. In particular, we focus on newcomers (first-time image posters). Is there evidence that AI-generated images make it easier for newcomers to contribute content – and if so, do these newcomers stay and contribute over time? Does this pattern diverge for those who are making Transparent AI posts to AI-disallowed versus AI-neutral communities?
Overall, the participation of newcomers was significantly more emphasized in making Transparent AI posts than the overall presence of newcomers in posting images to these communities. In AI-disallowed communities, newcomers usually publish 6.71% of all image-based posts. This ratio is similar in AI-neutral communities, with 6.51% of newcomers publishing image-based posts. However, newcomers made as much as 25.15% of Transparent AI posts in AI-disallowed communities during our period of analysis. In AI-neutral communities, newcomers made a similar 22.99% of the Transparent AI posts. The differences between newcomer production of identified AI content versus regular content was significant for both types of communities (Chi-square, $p<0.00001$).This over-representation in newcomers posting AI-generated content suggests that this kind of content may encourage and broaden community participation.
@@
Overall, we found a significant drop off for AI-posting newcomers in AI-disallowed communities compared to AI posters in AI-neutral communities as well as compared to newcomers posting non-AI content. Of Transparent AI newcomers in the AI-disallowed communities, only 8.6% remained active creators after a month. This rate was as high as 17.8% for Transparent AI newcomers in the AI-neutral communities [^13] This observed difference also suggests that our categorization of subreddits accurately reflects their tolerance towards generative AI.. The rate of Transparent AI newcomers remaining active for the AI-neutral communities was similar for the overall rate for all new contributors (18.7%). This rate of newcomers who continued to participate was even higher, 23.5%, for these general newcomers in the AI-disallowed communities – making the clear discouraging effect of the AI-disallowed communities on newcomers posting AI even more apparent. The differences between Transparent AI newcomers and Image-based newcomers were highly significant (Chi-square test, $p<0.00001$) for the AI-disallowed communities, but not statistically significant for AI-neutral communities.
We did a similar analysis looking at newcomers who are still participating after three or six month. The patterns and differences between the communities remained similar, though, of course, with lower rates of continued participation as more time passed across all categories. Overall, the findings suggest that in communities where the posting of generative AI content is not explicitly prohibited, newcomers to the subreddit continue to naturally participate over time.
The analysis above shows the low return rate of posters of AI-generated content, but we were also interested whether the returning newcomers continue to transparently post AI-generated content. In AI-disallowed communities, when Transparent AI newcomers returned after more than one month, 49.8% of their subsequent image-based posts were also Transparent AI posts. This number was similar in AI-neutral communities at 50.9%, though, as we have seen above, there were relatively more returning posters in this group. Note that these numbers likely represent lower bounds for the proportion of AI-generated content in these subsequent posts, since it is possible that subsequent image-based posts were also AI-generated images, but were published without explicit acknowledgment by the poster.
@@
Feedback, Interactions and Sentiment
We investigate the interactions – responses and comments on identified or suspected AI-generated content, and how they are different, in tone or dynamics, from other content or across community types.
First, we look at engagement levels with Transparent AI posts across different communities. We examined the engagement with image-based posts using the upvote ratio as a metric, which represents the proportion of upvotes relative to the total number of votes (both upvotes and downvotes) received by a post. For this analysis, we compared the distributions of upvote ratios in two distinct periods: (1) January-March 2022 and (2) January-March 2023. These periods were selected to capture the differences across an entire year, particularly given the release of significant text-to-image generation tools and the peak in activity observed towards the end of 2022, as discussed in previous sections.
Figure 2 illustrates the cumulative distribution of posts across different upvote ratio ranges (X-axes), and the corresponding cumulative number of posts (Y-axes). The left panel displays the distributions for all image-based posts in the analyzed subreddits, while the right panel focuses on Transparent AI posts. Additionally, the upper panel shows posts within AI-disallowed communities, whereas the lower panel presents posts from AI-neutral communities. The figure shows that the upvote ratio distribution for all image-based posts across both community types was similar (top left and bottom panels), and that there were only minor differences (dotted vs. solid line) between 2022 and 2023 upvote ratios in both.

### v2

@@ 4. Classification of the Dataset Using LLMs
Having gathered a dataset of potentially AI-relevant image-based posts based on keyword match, we further classified the posts using large language models (LLMs) (Abburi et al., 2023; Zhang et al., 2024; Pangakis and Wolken, 2024). First, since keyword matching can be imprecise, we use classifiers to filter the dataset to only retain posts and comments genuinely pertaining to AI-generated images. Second, we differentiate between text that identifies the image in the post as being AI-generated, versus those that broadly mention AI-generated images.
@@
To this end, we developed textual prompts and utilized the ChatGPT API [^5] https://platform.openai.com/docs/api-reference/introduction (GPT-4o model). Three distinct prompts were developed to correspond with each of the three type of text in our dataset: post’s title or body, comments made by the author of the post, and comments by other users (see Appendix C). These prompts were designed following the principles of chain-of-thought prompting (Wei et al., 2022). Each prompt required the language model to provide reasoning before giving a classification decision. Additionally, each prompt included nine examples from the dataset, implementing an in-context few-shot learning approach (Wei et al., 2022).
@@
We evaluated the accuracy of each of the three prompts using three validation sets. We randomly sampled 100 texts from each type, and manually labeled them according to the established coding. That is, the texts were labeled as follows: (A) Implies AI involvement, when the text explicitly or implicitly suggests that AI tools were involved in generating the attached image; (B) Does not imply AI, when the text mentions AI-generated images, but does not imply that the attached image is AI-generated (including texts that suggests that the attached image is not AI-generated); or (C) Unrelated to AI, only when the text is completely unrelated to AI-generated images (regardless of whether the text refers to the image in the post)—in most cases, a false positive of the keyword match. Next, we classified the three validation sets using the corresponding prompts. The prompts, along with examples of input texts and their corresponding classifications, are provided in Appendix C. This classification task achieved an accuracy of 93% for the title of the posts, 82% for comments made by the author of the post, and 84% for comments from other accounts. In comparison, the baseline accuracy for the majority class is 87% for post titles, 49% for comments made by the post’s author, and 55% for comments from other accounts. Note that the majority class differs across the three validation sets: for post titles and comments made by the post’s author, the majority label is Implies AI involvement, while for comments from other accounts, the majority label is Does not imply AI.
@@
We classified all data entities in the dataset using these prompts, following a process inspired by the QuaLLM framework (Rao et al., 2024). According to this process we submitted the prompt, consisting in each case of the relevant class explanation and the in-context examples, with every 100 input texts at a time [^6] Note that during the evaluation phase we also compared the accuracy values obtained by submitting different numbers of entities as input (i.e., 10 instead of 100) and achieved identical results.. Similarly, we requested that the LLM return the final output as a dictionary, including both the reasoning and classifications for all these 100 texts.
@@
In this paper, we focus mainly on texts, either in posts or comments, that imply that the image attached to the post was AI-generated (the first category of labels above). First, looking at people who make image posts in these communities, we identify on posts where the author implied or admitted that the image was AI-generated, whether in the title, or in one of the comments. We refer to posts where the creator itself implied the image was AI generated as Transparent AI posts. There are 5,998 distinct posts like these where the creator provides this information in the title (5,144 times) and/or comments (1,243). Second, we are interested in posts where other users or accounts make comments that refer in some way to the posted image as AI-generated. We refer to these posts where other accounts highlighted AI use as Suspected AI posts. There were a total of 7,322 comments from other accounts claiming AI was used, made on 4,375 distinct Suspected AI posts. The overlap between the two sets of posts was small—576 posts where the author transparently posted AI and commenters specifically referred to AI generation in their comments.
5. Categorization of the Communities
@@
We categorized the 57 subreddits into two types of communities: those with rules that explicitly ban AI images, and those whose rules do not mention AI. Previous papers have explored the reactions of community moderators to AI-generated content, including focusing on the decisions of Reddit moderators to enact rules restricting the use of AI, as well as the responses of their communities (Lloyd et al., 2025b; Shankar and Sim, 2024). We set out to distinguish the different communities based on their expressed approach to AI content. Using the data provided via the PRAW API [^7] https://praw.readthedocs.io/en/latest/index.html, we examined each subreddit’s API description, public description, and rules. If any of these fields indicated that AI-generated content (or particularly AI-generated images or art) is not permitted, we categorized the subreddit as AI-disallowed community. Otherwise, if the publication of AI-generated content was not mentioned in the subreddit’s description or rules, it was categorized as AI-neutral community [^8] In our dataset, no subreddit explicitly allows AI art in their description or rules.. Among the 57 subreddits, 24 explicitly forbid AI, while the remaining 33 subreddits do not reference AI in their descriptions and rules.
@@
To obtain a more comprehensive understanding and baseline for the subreddits in each category, we utilized torrents (Watchful1, 2024) to count the total number of posts that included a media item in these communities, regardless of whether they contained the AI-related keywords. We refer to these posts that included a media item, such as an image or video, as all image-based posts.
@@
Table 1. Characteristics of the AI-disallowed and AI-neutral communities.
This quantitative comparison reveals several unexpected findings. Although the majority of posts of any kind within these subreddits are image-based, comprising 75% in AI-disallowed communities and 77% in AI-neutral communities, only a minimal percentage of the posts included texts implying AI involvement in creating the image. Specifically, AI-claiming texts were presented only in 0.21% of image-based posts in AI-disallowed communities, and 0.12% in AI-neutral communities, including texts made by the author of the post (Transparent AI posts) and those made by commenters (Suspected AI posts). Since AIGC was not prevalent before 2022, we recalculated the metric using data from January 1, 2022, onward. Across 2022–2023, Transparent or Suspected AI posts accounted for only 0.49% of all image-based posts in both AI-disallowed and AI-neutral communities. Furthermore, when AI involvement was mentioned in relation to the creation of the image, it was mostly by the content creators themselves, accounting for 57% (4,458 out of 7,792) of cases in AI-disallowed communities, and 60% (1,540 out of 2,581) in AI-neutral communities. Finally, for 87% (2,894 out of 3,334) of the Suspected AI posts in the AI-disallowed communities, i.e. posts where a non-author commenter implied AI involvement, the comment served as the first indication that the attached image was AI-generated, without the author of the post disclosing the use of AI. The same percentage was observed in the second category as well. We explore these differences further in our analysis.
6. Analyses and Findings
@@
In this section, we present a comprehensive analysis to address our research questions. Using the dataset we collected and refined in previous sections, we examine variations in volume, participation, and feedback patterns across AI-disallowed communities and AI-neutral communities. This analysis highlights the distinctions between these two types of communities, as well as the differences in trends for the different types of posts — Transparent AI posts and Suspected AI posts.
6.1. Rise of AI Posts and Comments
@@
Figure 1. The volume of posts over time in AI-disallowed communities (red) and AI-neutral communities (green). The x-axis of the figure represents time (month), and the y-axis represents the number of posts. Plots 1-2 (top two) show the total volume of all image-based posts ; plots 3-4 show Transparent AI posts, and plots (5-6) show the number of Suspected AI posts.
@@
To address our first research question, we conducted an analysis of post volume over time across different communities to understand how posting patterns shifted in response to the introduction and increasingly mainstream adoption of generative AI throughout 2022. Specifically, we examined temporal patterns of posts containing AI-generated images according to our classification, and compared these trends to the overall volume of image-based posts within the subreddits we examined. We explore how these trends vary over time between the AI-disallowed communities and the AI-neutral communities. Figure 1 shows the number of posts over time between both types of communities: AI-disallowed communities in red, and AI-neutral communities in green. For brevity, the figure includes only posts published from the beginning of 2020 onward, as generative AI was not dominant prior to that period. The figure presents the number of posts for various data segments, enumerating from the top panel: (1-2) all image-based posts, (3-4) Transparent AI posts when authors admitted to posting AI-generated images, and (5-6) Suspected AI posts when commenters implied that the image in the post is AI-generated. For instance, Panel 2 shows that between March and August 2020, approximately 35,000 to 40,000 image-based posts were published in AI-neutral communities (which do not mention AI posting in their community rules). In contrast, two years later (March-August 2022), only approximately 25,000 to 30,000 image-based posts were published in the same subreddits.
@@
While the overall number of image-based posts has been declining since August 2021 in both community types (panels 1-2), the number of Transparent AI posts began to increase during the same period (panels 3-4). Though analyzing general trends in image-based posts is beyond the scope of this work, this difference suggests that the rise in generative AI posts is not due to broader trends in posting volume across subreddits, but rather specific trends related to AI-generated images.
@@
To check this, we calculated the timelines using the percentage values for each subreddit (instead of volume values) – an unbiased metric in terms of the varying volumes of subreddits – averaged for each community category. We revealed that the same trends persisted, indicating that the observed patterns were consistent and not dominated by the largest subreddits. Furthermore, we examined the 12 subreddits with the highest number of posts, and observed similar trends in the timelines, despite the differences in post volumes.
@@
6.2. Participation and Newcomers
The previous sections focused on the volume of posts and discussions over time and in relation to other image posts in the same communities. In this section, we address our second research question by examining participation patterns in the creation of AI-related posts across the different art communities. In particular, we focus on newcomers (first-time image posters). Is there evidence that AI-generated images make it easier for newcomers to contribute content – and if so, do these newcomers stay and contribute over time? Does this pattern diverge for those who are making Transparent AI posts to AI-disallowed versus AI-neutral communities?
Overall, the participation of newcomers was significantly more emphasized in making Transparent AI posts than the overall presence of newcomers in posting images to these communities. In AI-disallowed communities, newcomers usually publish 6.71% of all image-based posts. This ratio is similar in AI-neutral communities, with 6.51% of newcomers publishing image-based posts. However, newcomers made as much as 25.15% of Transparent AI posts in AI-disallowed communities during our period of analysis. In AI-neutral communities, newcomers made a similar 22.99% of the Transparent AI posts. The differences between the newcomers’ rate of production of identified AI content versus regular content was significant for both types of communities (Chi-square, $p<0.00001$). This over-representation in newcomers posting AI-generated content suggests that this kind of content may encourage and broaden community participation.
@@ 6.2. Participation and Newcomers
Overall, we found a significant drop off for AI-posting newcomers in AI-disallowed communities compared to AI posters in AI-neutral communities as well as compared to newcomers posting non-AI content. Of Transparent AI newcomers in the AI-disallowed communities, only 8.6% remained active creators after a month. In the AI-neutral communities, this rate of return was as high as 17.8% for Transparent AI newcomers. This observed difference also suggests that our categorization of subreddits accurately reflects their tolerance towards generative AI. The rate of Transparent AI newcomers remaining active for the AI-neutral communities was similar for the overall rate for all new contributors (18.7%) in these communities. This rate of newcomers who continued to participate was even higher, 23.5%, for these general newcomers in the AI-disallowed communities – making the clear discouraging effect of the AI-disallowed communities on newcomers posting AI even more apparent. The differences between Transparent AI newcomers and Image-based newcomers were highly significant (Chi-square test, $p<0.00001$) for the AI-disallowed communities, but not statistically significant for AI-neutral communities.
We did a similar analysis looking at newcomers who are still participating after three or six month. The patterns and differences between the communities remained similar, though, of course, with lower rates of continued participation as more time passed across all categories. Overall, the findings suggest that in communities where the posting of generative AI content is not explicitly prohibited, newcomers who use AI as an entry point for participation in the subreddit continue to naturally participate over time.
The analysis above shows the often-low return rate of posters of AI-generated content, but we were also interested in whether the returning newcomers continue to transparently post AI-generated content. In AI-disallowed communities, when Transparent AI newcomers returned after more than one month, 49.8% of their subsequent image-based posts were also Transparent AI posts. This number was similar in AI-neutral communities at 50.9%, though, as we have seen above, there were relatively more returning posters in this group. Note that all these numbers build on counting Transparent AI posts and thus are likely representing lower bounds for the proportion of actual AI-generated content in the initial and subsequent posts, since it is possible that AI-generated images are published without explicit acknowledgment by the poster.
@@
6.3. Feedback, Interactions and Sentiment
To address our third research question, we investigate the interactions—responses and comments—with identified or suspected AI-generated content, and how they are different in tone or dynamics from other content or across community types.
First, we look at engagement levels with Transparent AI posts across different communities. We examined the engagement with image-based posts using the upvote ratio as a metric, which represents the proportion of upvotes relative to the total number of votes (both upvotes and downvotes) received by a post. For this analysis, we compared the distributions of upvote ratios in two distinct periods: (1) January-March 2022 and (2) January-March 2023. These periods were selected to capture the differences over the span of a year, particularly given the release of significant text-to-image generation tools and the peak in activity observed towards the end of 2022, as discussed in previous sections.
Figure 2 illustrates the cumulative distribution of posts across different upvote ratio ranges (X-axes), and the corresponding cumulative number of posts (Y-axes). The left column panels display the distributions for all image-based posts in the analyzed subreddits, while the right panel focuses on Transparent AI posts. The upper row panels show posts within AI-disallowed communities, whereas the lower panels present posts from AI-neutral communities. The figure shows similar baseline response rates: the upvote ratio distribution for all image-based posts across both community types was similar (top left and bottom left panels), and there were only minor differences (dotted vs. solid line) between 2022 and 2023 upvote ratios in both.